In a city, the daily consumption of Electric power in millions of ‘kwh’ can be considered as a random variable following Gamma Distribution with α= 3, λ= ½. If the power plant in that city has a daily consumption of 12 million KWs, what is the probability that this supply of power is ‘insufficient’ in any given day?[ Hint: Find P(X>12)]
1
Expert's answer
2020-09-13T14:59:48-0400

We can find P(X>12) in a way

"P(X>12) = 1 - P(X \leq 12)" .

By definition, "P(X\leq x)" "=F(x)" - cumulative distribution function(CDF). For gamma distribution CDF is

"P(X\leq x) = F(x; \alpha, \lambda) = \frac{\gamma(\alpha, \lambda x)}{\Gamma(\alpha)}"

where where "{\displaystyle \gamma (\alpha ,\lambda x)}" is the lower incomplete gamma function and "\Gamma(\alpha)" is ordinary gamma function.

"P(X\leq12) = \frac{\gamma(3, 6)}{\Gamma(3)} = 0.938."

So, "P(X>12) = 1 -0.938 = 0.062"